OpenStudy (anonymous):

Every day, Marco runs for 45 minutes and covers 5 miles. How long would it take him to run 2 miles?

OpenStudy (compassionate):

Take 5 and divide it by 45, then multiply it by 2.

OpenStudy (anonymous):

0.2 repeating

OpenStudy (anonymous):

Suppose he runs at constant speed. Then it takes him 45/5 = 9 minutes to cover 5/5 = 1 mile. Doing that twice takes 2*9 = 18 minutes.